Card is drawn from an ordinary deck, Find the probability of getting a 3 or a diamond carf

1
Expert's answer
2022-03-09T13:30:27-0500

"P=\frac{13+4-1}{52}=\frac{16}{52}=\frac{4}{13}\approx 0.3077."